### v2

# PositionIC: Unified Position and Identity Consistency for Image Customization

Abstract
Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained entity-level spatial control remains elusive, hindering the broader real-world application. This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues. To this end, we introduce PositionIC, a unified framework that enforces position and identity consistency for multi-subject customization. We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence. On top of these data, we design a lightweight positional modulation layer that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity. Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task. PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
Figure 1: Rusults from PositionIC across various controllable image customization tasks.
Introduction
Recently, diffusion-based models have achieved remarkable advances, especially in text-to-image (T2I) generation (Podell et al. 2023; Cai et al. 2025; Rombach et al. 2022; Li et al. 2024, ; Labs 2024). These models can produce images with superior visual quality and high semantic fidelity given short or descriptive prompts.
Among the tasks of image generation, image customization generation has attracted considerable attention in recent years. It requires models to generate images that not only adhere to textual descriptions but also preserve the identity features of reference objects. Although recent methods (Wu et al. 2025; Mou et al. 2025; Tan et al. 2024; Wang et al. 2025) have achieved notable improvements in image fidelity, they still lack customized control over the position, size, and scale of generated subjects and suffer from the limitation of insufficient high-quality customized paired data. While other methods (Wang et al. 2024c; Li et al. 2023) permit the placement of specific objects at specific positions, their position controllability and consistency are still insufficient and underexplored. This limitation hinders the effectiveness in applications where precise spatial arrangement and accurate visual quality are demanded, such as e-commerce product display, storybook illustration, and interior arrangement. As a result, there remains pressing challenges for approaches that can simultaneously ensure high image fidelity and offer flexible, fine-grained control over image attributes.
In this paper, we propose an automatic data curation pipeline and PositionIC, a novel method for controllable image customization to tackle these two challenges, i.e., (1) lack of high-quality customized paired data with position signals; and (2) controllable subject-driven customization guided by the layout. Specifically, regarding the data challenge, we introduce a Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline. We utilize a hierarchical training strategy to build a bidirectional generation paradigm from single-object to multi-object and vice versa, which could lift resolution constraints. We adopt Multi-modal Large Language Models (MLLMs) on data filtering to improve data quality. However, MLLMs are prone to visual hallucinations (Han et al. 2024; Pi et al. 2024), which can lead to inaccuracies when comparing visual information. Instead of directly relying on MLLMs to assess the consistency between two images, we carefully design a data filter that leverages expert models to convert visual information into textual descriptions. These descriptions are then used to assist MLLMs in performing multi-dimensional consistency evaluation. For controllable image customization, we propose PositionIC, a lightweight position-controllable framework that decouples layout from subject. It unlocks the fine-grained position control capability of Diffusion Transformers (DiT) through instance- level attention. Additionally, PositionIC does not introduce any additional parameters thus no additional computation is required. By restricting reference image features to focus only on specific regions, PositionIC can generate customized subjects at desired locations, which greatly expands the applications of controllable image customization as shown in Figure 1. Further experiments demonstrate the outstanding universality and remarkable capability of our method.
Our contributions are summarized as follows:
  • We design an automatic framework for data synthesis to obtain high-fidelity paired data, addressing the lack of precise subject consistency and positional control signals in existing public datasets.
  • We propose PositionIC, a lightweight position-controllable framework that decouples layout from subject. It achieves precise placement of multiple subjects through regional attention. At the same time, it explicitly enhances the subject fidelity of Diffusion transformers.
  • Extensive experimental results demonstrate that our method not only achieves the state-of-the-art performances on image customization, but also exhibits the highest precision control capability.
Figure 2: Bidirectional Multi-dimensional Perception Data Synthesis framework. (a) We use Subject200K to train a weak model. (b) Forward generation of multi-subject data pairs. (c) Reverse generation of multi-subject data pairs. (d) We utilize MLLMs to filter out our data pairs.
Figure 3: The overall framework of PositionIC. (a) Reference images and prompts are encoded and concatenated with the latent embeddings $z_{t}$, then the whole token sequence is passed to the double stream blocks of DiT. (b) The mechanism of attention horizon. Each reference image $z_{r}^{i}$ is only visible for the specific area of latent noise $z_{t}$ in the attention map.
Related Work
Subject-driven Generation
In addition to using text prompts for conditional image generation, current diffusion models (Wei et al. 2023; Zhang et al. 2024; Huang et al. 2024; Xiao et al. 2025; Huang et al. 2024) support reference image input to achieve preservation of subject identity. Dreambooth (Ruiz et al. 2023) and LoRa (Hu et al. 2022) control the generation of diffusion models through fine-tuning on the specific subject. IP adapter (Ye et al. 2023) takes the reference images as feature inputs and achieves image-to-image generation by injecting the feature into the cross-attention module of the source model. Recently, DiT-based subject-driven models (Wu et al. 2025; Tan et al. 2024, 2025; Mou et al. 2025; Labs et al. 2025) further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation process.
Position Controllable Generation
Some works (Wang et al. 2024b; Chen et al. 2024; Wang et al. 2024a; Jiménez 2023; Bar-Tal et al. 2023; Shi et al. 2025) have attempted to generate with precise layout control. Gligen (Li et al. 2023) encodes Fourier embedding as grounding tokens to inject position information. MS-diffusion (Wang et al. 2024c) utilizes a grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinder further application.
Method
Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
Customized Data Paired Synthesis
We divide the automated data generation process into three stages. (1) Inspired by UNO, we first train a weak model for image customization tasks using Subject200K (Tan et al. 2024) dataset. The generated results are then segmented and passed through a Flux-Outpainting model with random placement to inject spatial control information. The paired data obtained in this stage is filtered by a filter to ensure high fidelity. These data are then used to train the PositionIC-Single model. (2) In the second stage, we perform forward generation of multi-image data pairs. We independently process Subject200K samples via PositionIC-Single, then randomly pair and position the output as input to the Flux-Outpainting model, thereby obtaining multi-subject paired data. (3) To enhance data diversity and improve generalization, we reverse the above process in the third stage. We first use Multi-modal Large Language Models (MLLMs) to generate text descriptions containing multiple subjects, then employ the Flux model to create high-resolution images. Objects are detected and cropped from these results, individually processed by the PositionIC-Single, resulting in high-resolution multi-subject paired data.
Through bidirectional data synthesis, we construct PIC-400K dataset with noise interference.
Figure 4: Qualitative comparison of single-subject generation with different methods on DreamBench.
Figure 5: Qualitative comparison of multi-subject generation with different methods on DreamBench. We adopt a fixed bounding box (e.g., bottom left and bottom right) for generation.
Figure 6: Inaccuracy of directly using CLIP-I and DINO. The revised score is less affected by the size of the subject which can reflect the subject consistency more authentically.
Multi-dimensional Perception Data Filter
Previous studies have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color, and details. Lastly, we employ Describe Anything Model (DAM) (Lian et al. 2025), a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
Then we calculate the ranking separately and average them for each image pair to obtain final ranks:
$rank=avg(r(s_{v}),r(s_{vlm}),r(s_{ds})),$ (1)
where $r(\cdot)$ denotes rank. With a lower rank indicating a higher subject similarity.
We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs.
Regional Attention Horizon
To manage subject consistency generation position simultaneously (e.g., generating the subject in a specific location in the image), we introduce a light-weight approach named attention accumulating, which can unlock DiT’s ability of spatial control without extra training and inference cost. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text attention. If extending to single or multi-subject custom generation via concatenating method like (Wu et al. 2025; Tan et al. 2024; Labs et al. 2025; Zhang et al. 2025), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulating, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon. Previous works (Zhang et al. 2025; Chen et al. 2024) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
As shown in Figure 3, we first encode the reference images $I_{r}^{i}$ to the latent space via VAE $\epsilon(\cdot)$. Encoded results $z_{ref}$ are then concatenated with noise latents $z_{t}$ and text embedding $z_{p}$. This process can be formulated as:
$z_{ref}=[\epsilon(I_{r}^{1}),\epsilon(I_{r}^{2}),...,\epsilon(I_{r}^{N})],$
$z=Concatenate(z_{p},z_{t},z_{ref}),$ (2)
where $N$ is the number of reference images and $z$ denotes the input tokens to the DiT model.
Since the global text prompt contains information about the entire image, generating a reference object in a specific region requires focused and singular attention. They only need to pay attention to the corresponding region while ignoring other irrelevant tokens. Therefore, a limited attention horizon is required. We use a binary attention mask to shield the regions that each token should not directly focus on. As shown in Figure 3, each reference image $z_{ref}$ has a restricted attention horizon, which blocks the attention between itself and other reference images. Moreover, only the area within the bounding box in the noise is visible for the specific reference image. The mechanism of attention horizon mask $M$ can be formulated as:
$M(z_{ref}^{i},z_{ref}^{j})=0,i\neq j,$
$M(z_{ref}^{i},z_{t}^{n})=0,z_{t}^{n}\notin BOX_{i},$
$M(other)=1,$ (3)
where $BOX_{i}$ is the bounding box of $i^{th}$ reference subject in the noise and $z_{t}^{n}$ is the $n^{th}$ patch of noise. Thus, the computation of attention is derived as:
$Attention=Softmax(\frac{QK^{T}}{\sqrt{d}}+\log{M})\cdot V$ (4)
PositionIC-Bench
Most existing customized image evaluation (e.g., DreamBench) lacks explicit spatial position annotations. Thus, there is no universal data benchmark for evaluating subject-driven methods with position control. To address this gap, we propose PositionIC-Bench, a benchmark to evaluate subject consistency and position accuracy simultaneously.
We manually select 252 single-subject samples and 296 multi-subject samples in the benchmark, where the object bounding boxes conform to standard proportions and include challenging positional relationships.
To evaluate our model’s performance on the benchmark, We leverage multi-modal large language models (MLLMs) to detect objects and assess their content, computing the IoU and object category alignment between the generated and reference subjects.
Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench. RPF denotes Regional Prompting Flux. MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
Experiments
Implementations
Training Detail
Following UNO, we first initialize the model using FLUX.1 dev and apply UnoPE to extend the position embedding of the reference images to the non-overlapping area diagonally. We train a LoRA at the rank of 512 on 8 NVIDIA A100 GPUs and set the total batch size of 128. The learning rate is set to ${10}^{-5}$ with cosine warm up. In the first stage, we train the single-subject model on 44k single-subject pairs for 10k steps. We then continue our training on 54k multi-subject pairs for 8k steps, which extends the multi-subject generation capability to the model obtained in the first stage.
[…]
Dreambooth 0.776 0.215 0.679
BLIP-Diffusion 0.787 0.234 0.742
ELITE 0.803 0.235 0.723
SSR-Encoder 0.797 0.206 0.725
RealCustom 0.783 0.242 0.765
MS-Diffusion 0.808 0.242 0.791
OminiControl 0.805 0.268 0.735
OmniGen 0.791 0.267 0.751
UNO 0.840 0.253 0.814
DreamO 0.835 0.258 0.802
PositionIC$({Ours})$ 0.846 0.269 0.823
Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the best and the underlined value is the second best.
[…]
BLIP-Diffusion 0.703 0.212 0.541
MIP-Adapter 0.752 0.254 0.657
MS-Diffusion 0.772 0.261 0.683
OmniGen 0.749 0.291 0.668
UNO 0.781 0.279 0.707
DreamO 0.779 0.273 0.698
PositionIC$({Ours})$ 0.819 0.279 0.771
Table 2: Evaluation on DreamBench for multi-subject driven generation.
Evaluation Metrics
For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench (Ruiz et al. 2023). For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench We use Vision-R1 (Huang et al. 2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
Method Single-Subject Multi-Subject
[…]
Gligen (Li et al. 2023) 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
RPF (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
MS-Diffusion 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
PositionIC$({Ours})$ 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
Table 3: Quantitative results of controllable spacial generation on PositionIC-Bench.
Qualitative Result
We visualize the comparison results with the current state-of-the-art methods at Figure 4. Overall, our method surpasses all current methods in terms of visual effects. It can be seen that our method can still effectively follow the prompt while maintaining subject consistency, demonstrating higher text fidelity. Other methods either fail to follow complex instructions or cannot maintain consistency. (e.g., UNO and DreamO cannot reproduce the dog face, while SSR-Encoder fails to add the Santa hat.) The results of the third row also reveal that PositionIC has a great capability on patterns and text. Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility.
Figure 5 shows the comparison of multi-subject generation. To control variables, the images generated by our method adopt a fixed bounding box (e.g., bottom left and bottom right). It can be seen that in the more difficult multi-subject scenario, PositionIC can still maintain high subject similarity and follow the given text prompt, whereas results from most other methods fail to preserve consistency for each subject and even ignore certain subjects.
We further evaluate the positional control capability of existing methods on PositionIC-Bench via randomly generating the bounding boxes. As shown in Figure 7, PositionIC accurately generates subjects that fully occupy the bounding box without damaging their features. More importantly, our flux-based approach significantly outperforms others in terms of image aesthetic quality and the logical coherence of multi-subject compositions.
Figure 8: Ablation study of data filter. We train our model on PIC-98K, PIC-400K and Subject200K respectively.
Quantitative Evaluations
Subject-driven Analyses
Specifically, we discover that different object sizes can lead to inaccurate scores. As shown in Figure 6, . To avoid the sensitivity, we crop the subject from original image as source images for evaluation.
We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject. As presented in Table 1 and Table 2, PositionIC achieves the highest scores on both CLIP-I and DINO of 0.846 and 0.823 on single-subject, 0.819 and 0.771 on multi-subject, respectively. PositionIC also has competitive CLIP-T compared to existing methods. The evaluation results indicate that PositionIC has remarkable performance on subject consistency and text fidelity.
Controllable Spacial Generation
Table 3 presents the results of spatial control evaluation. PositionIC achieves superior performance in both single-subject and multi-subject position control. For single-subject position control, PositionIC has the highest IoU of 0.828 across all methods and has competitive $AP$ scores compared with Gligen. For multi-subject evaluation, PositionIC achieves the highest scores in both AP and IoU scores, demonstrating a significant advantage over existing methods.
Ablation Study
In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 8 and Table 4.
Impact of data filter. Results in Figure 8 illustrate the advanced fidelity of BMPDS. We directly trained with PIC-98K, PIC-400K and the Subject200K dataset without injecting position control information. PIC-400K achieves significantly higher CLIP-I scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
[…]
w/o horizon mask 0.784 0.269 0.686
w/ horizon mask 0.846 0.269 0.823
Table 4: Ablation study of horizon mask. Our model perform better subject fidelity on Dreambench after restrict the attention horizon of reference images.
Horizon mask. As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask. We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistentcy.
Conclusion
In this work, we present PositionIC, an innovative framework capable of customizing multiple subjects with precise position control. PositionIC decouples layout signal from subject feature without introducing additional parameters and training cost. Additionally, we carefully design an automatic data curation framework to obtain high-fidelity paired data. We adopt bidirectional generation and present a multidimensional perception filter to improve object consistency in acquired data. Extensive experiments demonstrate that PositionIC performs high-quality generation in both single-subject and multi-subject consistency, as well as in controllable subject positioning. We hope our work can advance the development of controllable image customization.